### v1

# Coherence-Assisted Entanglement Activation During Pure Dephasing Evolution

Abstract
We investigate the system-environment entanglement activation during pure dephasing evolution, also called controlled unitary. A lower bound of generated entanglement is formulated with the initial coherence in the system and using relative entropy measure. For a qubit-environment case, this lower becomes a closed form by replacing conditional entropy with the overlap on the states in environment. Applying this result to the spin-boson model, we illustrate how the lower bound evolves as a function of time multiplied cutoff frequency, shown plots. The different behaviors of lower bounds with respect to Ohmicity, s=2 and s=3, corresponding to Markovian and non-Markovian evolution. In addition, we show that the decoherence in the system throughout the pure dephasing evolution represents another lower bound and upper bound of generated entanglement, indicating that larger entanglement generation leads more decoherence.
1 Introduction
Entanglement Horodecki et al. (2009) has been the subject of quite amount of researches in last decades due to its significance not only for quantum information and computation, but for the foundations of quantum mechanics as well. Understandably, the vast majority of entanglement studies focus on the purposeful entanglement generation, manipulation, and preservation. However, there are processes where entanglement can be generated spontaneously. Open systems dynamics (see e.g. Schlosshauer (2007); Breuer and Petruccione (2002) for a modern introduction) is one of such examples, where the interaction between the system and its environment leads in general to entanglement creation.
Until recently Roszak and Cywiński (2015); Roszak (2018, 2020); Rzepkowski and Roszak (2020) this system-environment entanglement has not been systematically studied using modern methods of quantum information theory, being mostly treated as a mere source of decoherence. However, with the emergence and the gradual acceptance of a new paradigm in open systems, given by the quantum Darwinism idea Zurek (2009) and its refinements such as Spectrum Broadcast Structures and strong Darwinism (see Korbicz (2021) for a review), the environment has been recognized as a carrier of valuable information about the system. Therefore, studying system-environment entanglement is important not only for understanding decoherence processes, since entanglement in general increases decoherence efficiency, see e.g. Takahashi et al. (2022) for recent take on this question, but also to know how environment acquires system-related information.
One can illustrate system-environment entanglement generation on a paradigmatic example of the, so called, controlled unitary evolution, alternatively called also a measurement-type or a pure dephasing evolution, and defined by:
$|s_{i}\rangle|E_{0}\rangle\rightarrow|s_{i}\rangle|E_{i}\rangle,$ (1)
in some preferred basis of the system $|s_{i}\rangle$, called the pointer basis  Zurek (1981, 1982, 2003). This basis is specified by the nature of the interaction, e.g. the dipole approximation for the electromagnetic interaction, leading to the spin-boson model Schlosshauer (2007); Breuer and Petruccione (2002). Whenever there is a superposition present in the initial state $|\psi_{S}\rangle=\sum_{i}a_{i}|s_{i}\rangle$, the resulting state $|\psi^{\prime}_{SE}\rangle$ is entangled:
$|\psi_{S}\rangle|E_{0}\rangle\rightarrow|\psi^{\prime}_{SE}\rangle=\sum_{i}a_{i}|s_{i}\rangle|E_{i}\rangle$ (2)
while the reduced subsystem $\textrm{Tr}_{E}[|\psi^{\prime}_{SE}\rangle\langle\psi^{\prime}_{SE}|]$ lost the initial coherence in $|\psi_{S}\rangle$. In the modern language of the resource theories Chitambar and Gour (2019), this an example of the resource interconversion, where one resource, coherence in the pointer basis, gets interconverted into another - entanglement  Piani et al. (2011); Chuan et al. (2012); Streltsov et al. (2015); Killoran et al. (2016); Qiao et al. (2018); Regula et al. (2018). Let us remark here that an abstract coherence theory finds its natural application in open systems with pure dephasing evolution as the latter contains a natural, preferred basis.
However easy the pure state interconversion (2) looks like, the situation gets much more complicated when considering mixed states, especially if one wants to quantify the amount of the generated entanglement. Assuming that the initial state of the system is pure, Roszak et al.  Roszak and Cywiński (2015); Roszak (2018) recently found necessary and sufficient separability conditions for the system-environment split and have developed own operational measure of the system-environment entanglement Roszak (2020); Rzepkowski and Roszak (2020). However, it is still unknown how to generalize their methods to mixed initial states of the system.
Previous studies of the conversion from coherence to entanglement showed that the entanglement generated by incoherent operations between coherent state in the system and incoherent pure state in environment is upper bounded by the initial coherence in the system respect to relative entropy measure Streltsov et al. (2015). However, for given Hamiltonian, the corresponding pure dephasing evolution is not incoherent operation generally and we also take the environmental state as arbitrary, their finding does not hold in our case. In pure dephasing evolution, therefore, it is not clear how entanglement generation is quantified through the initial coherence or decoherence in the system.
In this work, we study entanglement generation for mixed states of the system. We derive two lower bounds and one upper bound for the system-environment entanglement as measured by the relative entropy of entanglement $E_{r}$ Vedral et al. (1997); Vedral and Plenio (1998). Following the simple intuition (2) and its developments Piani et al. (2011); Chuan et al. (2012); Streltsov et al. (2015); Killoran et al. (2016); Qiao et al. (2018); Regula et al. (2018), the first lower bound is given in terms of the initial coherence of the system with respect to the natural pointer basis defined by the pure dephasing evolution, diminished by the remaining uncertainty about the system after the best measurement on the environment.
The second lower bound and the upper bound are quantified with the decoherence in the system, following the intuition that entanglement generation requires not only the initial coherence but the decoherence the system after the evolution which is also suggested by (2).
Regarding the first lower bound, for two-dimensional central systems, i.e. for qubits, we derive a stronger result in given by a closed formula that links entanglement with the distinguishability of the environment states. In particular, this allows to study entanglement dynamics for open qubits, understood as the real-time monitoring of the minimum amount of entanglement produced during the evolution. As an illustration, we apply our results to the spin-boson model (see e.g. Schlosshauer (2007) and the references therein). We study the dependence of the minimum generated entanglement on various model parameters e.g. the temperature and Ohmicity. Varying the latter through the well-known Markovian/non-Markovian transition, we see the traces of non-Markovianity in the dynamics of entanglement; cf. Rivas et al. (2010).
2 Preliminaries
Pure dephasing evolution is defined by the following Hamiltonian:
$H=H_{S}+H_{E}+\sum_{i=0}^{N-1}|i\rangle\langle i|\otimes V_{i}.$ (3)
$H_{S}$ is the self Hamiltonian of the system which here we assume to be commuting with the interaction, $H_{S}=\sum_{i}\varepsilon_{i}|i\rangle\langle i|$. The case of a more general $H_{S}$ is out of the scope of this work as we concentrate on the pure dephasing models. $H_{E}$ and $V_{i}$’s are arbitrary environmental observables. (3) is the most general Hamiltonian that generates evolution of the type (2). The unitary evolution $U$ is given by:
$U(t)=\sum_{i}|i\rangle\langle i|\otimes U_{E}^{i}(t),\quad U_{E}^{i}(t)=e^{-(\varepsilon_{i}+H_{E}+V_{i})t},$ (4)
which is the controlled-unitary or measurement-type operation in the quantum information terminology. The basis $|i\rangle$ is the pointer basis. Since observables $V_{i}$ are arbitrary, in general non-commuting, there does not exist a common diagonalizing basis, i.e. an incoherent basis for the environment, which makes the whole unitary operator incoherent. We will review the definition of incoherent operations below.
We will assume that the system and the environment are initially product, i.e. $\rho_{SE}(0)=\rho_{S}\otimes\rho_{E}$. This assumption is justified here by the fact that we are interested in the build-up of correlations, thus we want to initiate the system in a completely uncorrelated state, so that any produced correlations are due to the interaction. Applying the unitary operator defined in (4), we obtain the following state $\sigma_{SE}$ after the interaction:
$\sigma_{SE}=U(\rho_{S}\otimes\rho_{E})U^{\dagger}=\sum_{i,j}|i\rangle\langle i|\rho_{S}|j\rangle\langle j|\otimes U^{i}_{E}\rho_{E}{U^{j}_{E}}^{\dagger}$ (5)
In order to evaluate the amount of $SE$ entanglement, we will use the relative entropy of entanglement $E_{r}$ Vedral et al. (1997); Vedral and Plenio (1998), defined as:
$E_{r}(\rho_{SE})=\min_{\xi_{SE}\in Sep}S(\rho_{SE}||\xi_{SE})$ (6)
where $S(\rho||\xi)=\textrm{Tr}[\rho(\log\rho-\log\xi)]$ is the quantum relative entropy and the minimization is taken over all separable states $\xi_{SE}=\sum_{i}p_{i}\xi_{S}^{i}\otimes\xi_{E}^{i}$. This is a proper measure of entanglement Plenio and Virmani (2007), with a clear operational interpretation as the exponential bound on asymptotic probability of error of confusing $\rho_{AB}$ with the nearest separable state via POVM measurements, see e.g. Vedral (2002). The minimization in definition is the reason why it is difficult to evaluate the value directly. We employ the following lemma to bypass this problem:
Lemma 1.
Plenio et al. (2000) The relative entropy of entanglement is bounded from below by the negative conditional entropy, i.e. for any bipartite state $\rho_{AB}$ we have:
$E_{r}(\rho_{AB})\geq\max\{S(\rho_{A})-S(\rho_{AB}),S(\rho_{B})-S(\rho_{AB})\}$ (7)
where $S(\rho)$ is the von Neumann entropy: $S(\rho)=-\textrm{Tr}[\rho\log\rho]$.
We will also need elements of the quantum coherence theory; see e.g.  Baumgratz et al. (2014). Since quantum coherence is a basis dependent resource, we need to specify which basis is incoherent. In principle, one can develop a theory with any incoherent basis. However, since we study a concrete family of open dynamics, given by (3), the pointer basis of the central system $|i\rangle$ is the natural candidate here because if the system loses coherence due to the interaction, decoheres to this basis. We do not specify the incoherent basis for environment because the coherence in the environment does not matter in our analysis as we will show below. In the language of the coherence theory, an incoherent state is simply a diagonal state in the incoherent basis, which is of course nothing else but a decohered state in the open systems/quantum foundations terminology. We denote by $\mathcal{I}$ the set of all incoherent states. In a resource theory, of which coherence theory is an example, an operation which does not create resource is called a free operation respect to that resource. Various free operations for coherence are proposed for different physical motivations Chitambar and Gour (2016). The most standard one is incoherent operation whose Kraus operators $\{K_{n}\}$ satisfy $\frac{K_{n}\rho K_{n}^{\dagger}}{\textrm{Tr}[K_{n}\rho K_{n}^{\dagger}]}\in\mathcal{I}$ for all $n$ and any incoherent state $\rho\in\mathcal{I}$. One can now define the relative entropy of coherence $C_{r}$ in analogy to the relative entropy of entanglement (6) Baumgratz et al. (2014):
$C_{r}(\rho)=\min_{\sigma\in\mathcal{I}}S(\rho||\sigma)=S(\Delta(\rho))-S(\rho)$ (8)
where $\Delta(\rho)$ is called the dephasing channel, defined as $\Delta(\rho)=\sum_{i}|i\rangle\langle i|\rho|i\rangle\langle i|$.
Another ingredient that we will need is a measure of state distinguishability. Out of the possible measures Fuchs and van de Graaf (1999), we will use the state fidelity $B(\rho,\sigma)$, also called generalized overlap, defined as:
$B(\rho,\sigma)=$ $\min_{\textrm{POVM}:\{M_{i}\}}\sum_{i}\sqrt{\textrm{Tr}[\rho M_{i}]\textrm{Tr}[\sigma M_{i}]}=\textrm{Tr}\sqrt{\rho^{\frac{1}{2}}\sigma\rho^{\frac{1}{2}}}$ (9)
where $\textrm{Tr}[\rho M_{i}]\textrm{Tr}[\sigma M_{i}]$ is the probability that one obtains the same value $i$ from the measurement on $\rho$ and $\sigma$. Vanishing state fidelity is equivalent to a perfect distinguishability of the two states by POVM’s. This justifies the use of the state fidelity as a measure of distinguishability. This measure does not have as clear operational meaning as e.g. the quantum Chernoff bound Audenaert et al. (2007), related similarly to the quantum relative entropy to the asymptotic discrimination error of two states, but it is much easier to work with due to the compact form (9) and it also provides both an upper and a lower bound for the quantum Chernoff information.
Discriminating states from an ensemble =$\mathcal{E}=\{p_{i},\rho_{i}\}$ is a more difficult task and its efficiency can be quantified by the, so called, accessible information $I_{acc}(\mathcal{E})$. It is defined as the maximum of the Shannon (classical) mutual information $I(i:m)$ Cover and Thomas (2006) between the ensemble index $i$ and the measurement result $m$, taken over all POVM’s $\{M_{m}\}$ on $\rho=\sum_{i}p_{i}\rho_{i}$. The formal expression is given by:
$I_{acc}=\max_{POVM:\{M_{m}\}}I(i:m),$ (10)
where $I(i:m)$ is defined using the conditional probabilities of measuring $m$ while the sate was $\rho_{i}$: $p(m|i)=\textrm{Tr}(M_{m}\rho_{i})$.
For two-element ensembles, the two notions of discrimination are connected by the following Lemma, which will be central for our reasoning:
Lemma 2.
Rahul and Ashwin (2007)[Lemma 3.6] Let =$\mathcal{E}=\{p,\rho_{0},(1-p),\rho_{1}\}$ be a two-element ensemble of quantum states of arbitrary finite dimensions. Then:
$I_{acc}(\mathcal{E})\geq h(p)-2\sqrt{p(1-p)}B(\rho_{0},\rho_{1})$ (11)
where $h$ is the binary Shannon entropy: $h(p)=-p\log_{2}p-(1-p)\log_{2}(1-p)$.
3 Bounds on the system-environment entanglement generation
In the pure dephasing, open evolution, described in the Introduction, the system’s information is transferred to the environment through the controlled-unitary interaction (4). This interaction can be interpreted as the environment implementing a measurement in the pointer basis $\{|i\rangle\}$ and storing the information in its state. Indeed, as one can easily prove, the state of the environment in the course of the evolution becomes:
$\sigma_{E}(t)=\textrm{Tr}_{S}[U(t)\rho_{S}\otimes\rho_{E}U(t)^{\dagger}]=\sum_{i}p_{i}\rho_{i}(t),$ (12)
where $p_{i}=\langle i|\rho_{S}|i\rangle$ are the initial pointer probabilities and the conditional states of the environment $\rho_{i}(t)=U_{E}^{i}(t)\rho_{E}{U_{E}^{i}}(t)^{\dagger}$ encode, in a generally imperfect way, the information about the state of the central system. One can now try to recover this information by implementing a POVM measurement $\{M_{m}\}$ on the environment in order to extract the value of $i$, encoded by the environment in the course of the evolution. One may view this as an indirect measurement of the system through the environment, common in the macroscopic world; cf. the studies of objectivity in quantum mechanics in e.g. Zurek (2009); Korbicz (2021). The uncertainty of this indirect measurement is represented by Shannnon conditional entropy $H(i|m)$ Cover and Thomas (2006). From above observations, it is natural to think the less this uncertainty is, the stronger the system and the environment are correlated. On the other hand, the more is there initial coherence in the pointer basis $\{|i\rangle\}$, the more entanglement should be generated Killoran et al. (2016). Indeed this is so as the following Lemma shows:
Lemma 3.
For any POVM measurement $\{M_{m}\}$ on the environment, the relative entropy of entanglement, generated by the pure dephasing evolution $U$, is bounded from below by:
$E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\geq C_{r}(\rho_{S})-H(i|m).$ (13)
Proof.
Applying Lemma 1 to $\sigma_{SE}=U\rho_{S}\otimes\rho_{E}U^{\dagger}$, we have
[…]
$S(\sigma_{E})-S(\sigma_{SE})\}.$ (14)
We focus on finding a lower bound for $S(\sigma_{E})-S(\sigma_{SE})$ because $S(\sigma_{E})$ can be transformed to the conditional entropy or the overlap of environmental states through lemma 2, leading us to a nice result represented in corollary 1. We will also consider a lower bound for $S(\sigma_{S})-S(\sigma_{SE})$ later. We first observe that:
$S(\sigma_{E})-S(\sigma_{SE})=S(\sigma_{E})-S(\rho_{S})-S(\rho_{E})$ (15)
where we used fact that the unitary evolution preserves the von Neumann entropy and it is additive for product states. Notice (15) can be negative while the relative entanglement should be always non-negative.
Next, let us recall the Holevo bound on the $I_{acc}$ associated with (12):
$I_{acc}\leq S(\sigma_{E})-\sum_{i}\langle i|\rho_{S}|i\rangle S(U_{i}\rho_{E}U_{i}^{\dagger}),$ (16)
from which it follows that:
$S(\sigma_{E})\geq I_{acc}(\sigma_{E})+S(\rho_{E}),$ (17)
since $S(U_{i}\rho_{E}U_{i}^{\dagger})=S(\rho_{E})$. From the definition of the Shannon mutual information $I(x:y)=H(x)-H(x|y)$, where $H(x)$ and $H(x|y)$ are the classical Shannon and relative entropies respectively, the accessible information $I_{acc}$ can be represented as follows.
$I_{acc}=\max_{POVM:\{M_{m}\}}I(i:m)=\max_{POVM:\{M_{m}\}}\left[H(i)-H(i|m)\right].$ (18)
Index $i$ is a random variable, distributed with the initial pointer probabilities $p_{i}$, as follows from (12), or equivalently with the eigenvalues of the fully decohered state $\Delta(\rho_{S})=\sum_{i}|i\rangle\langle i|\rho_{S}|i\rangle\langle i|$. Therefore we may write $H(i)=S(\Delta(\rho_{S}))$. Since there is a maximization over all POVM in the definition (18), for an arbitrary POVM $\{M_{m}\}$ we will have:
$I_{acc}\geq H(i)-H(i|m)=S(\Delta(\rho_{S}))-H(i|m)$ (19)
Now we recover the statement by substituting (17) and (19) into (15):
$E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\geq$ $S(\sigma_{E})-S(\rho_{S})-S(\rho_{E})$
$\geq$ $I_{acc}-S(\rho_{S})$ (20)
$\geq$ $S(\Delta(\rho_{S}))-H(i|m)-S(\rho_{S})$
$=$ $C_{r}(\rho_{S})-H(i|m),$
where in the final line we used the fact (8). ∎
The above Lemma is valid for any finete dimensionality of the central system and the environment. The price to pay for this generality is the presence of the arbitrary POVM in the bound (15). This arbitrariness can be removed when the central system is a qubit. Then, there ensemble (12) consists of only two states $\rho_{0}(t)=U_{0}(t)\rho_{E}U_{0}(t)^{\dagger}$, $\rho_{1}(t)=U_{1}(t)\rho_{E}U_{1}(t)^{\dagger}$ with probabilities $p=\langle 0|\rho_{S}|0\rangle$ and $1-p=\langle 1|\rho_{S}|1\rangle$. We can then apply the Lemma 2 and obtain:
Corollary 1.
If the central system is two-dimensional, then:
$E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})$
$\geq$ $C_{r}(\rho_{S})-2\sqrt{p(1-p)}B(\rho_{0}(t),\rho_{1}(t)).$ (21)
The proof follows immediately from the application of Lemma 2 in (20).
Let us look closely at Lemma 3 and give some interpretations. The second term $H(i|m)$ on the right hand side is the uncertainty about the system’s state measured by the environment after our best measurement on environment. Non zero $H(i|m)$ means that we cannot conclude the system’s value in the pointer basis even if we could use any POVM on the environment. Reversely, if $H(i|m)$ becomes zero, the information about the system is transferred fully to environment and entanglement must be there unless system’s coherence is zero, which means the system is fully classical state. After all, when the system has more coherence in pointer bases, more entanglement is required to know the value of system certainly. One can also re-arrange the inequality (13) and obtain the lower bound on the uncertainty about the system’s state in terms of initial coherences and the generated entanglement:
$\max_{POVM:\{M_{m}\}}H(i|m)\geq C_{r}(\rho_{S})-E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger}).$ (22)
Similar interpretation can be applied to the qubit version (1). If $p(1-p)=0$ one knows the system’s state with certainty. Notice that $C_{r}(\rho_{S})=0$ in this case. Also if $B(\rho_{0}(t),\rho_{1}(t))=0$, it implies that $\rho_{0}(t)$, $\rho_{1}(t)$ are perfectly distinguishable, which means maximum information has been transferred to the environment.
Note that the bounds (13),(1) are not tight and can even become negative as we have mentioned in the proof and will see in an example in the next section. In that case of course the bounds trivialize as $0\leq E_{r}$ always. Nevertheless, despite the lack of optimality, they provide a valuable information about the open evolution, certifying that at least a given amount of entanglement has been generated. This is particularly interesting when the right hand sides of (13),(1) are large, which implies a large amount of created entanglement.
Finally, we will give an upper bound and another lower bound which we mentioned in the proof of lemma 3 on the generated entanglement. From a previous research Streltsov et al. (2015), an upper bound on entanglement generated by incoherent operations with initially coherent system $\rho_{S}$ and an incoherent pure ancilla $|0\rangle\langle 0|_{E}$ is given by the initial coherence in system, for any measure of entanglement $E_{D}$ and coherence $C_{D}$ corresponding to a contractive distance $D$:
$E_{D}(\Lambda[\rho_{S}\otimes|0\rangle\langle 0|_{E}])\leq C_{D}(\rho_{S})$ (23)
where $\Lambda$ is any incoherent operation on the bipartite system. When we specify $D$ as the relative entropy, which we have been using throughout this work, the above upper bound becomes tight.
However, since the corresponding unitary evolution $U$ for given Hamiltonian of (3) is not incoherent in general regardless of incoherent basis for environment and we assume the environmental state arbitrary, the above upper bound is not valid in our scenario. Also, considering our intuitive picture of system-environment interaction where coherence is converted into entanglement, it is interesting to connect the generated entanglement not only to the initial coherence $C_{r}(\rho_{S})$ as we did in lemma 3 for lower bound, but to the decoherence, the difference of coherence between the initial and final states in system as we demonstrate below.
Lemma 4.
Let $D_{S}$ be the decoherence in the system defined as $D_{S}=C_{r}(\rho_{S})-C_{r}(\sigma_{S})$. Then the relative entropy of entanglement generated by pure dephasing evolution $U$ is bounded by below and above as:
$D_{S}-S(\rho_{E})\leq E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\leq D_{S}+H(i).$ (24)
Proof.
Let us prove the left inequality first. From lemma 1, $E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\geq S(\sigma_{S})-S(\sigma_{SE})$. The invariancy of entropy over unitary evolution and the fact $S(\Delta(\sigma_{S}))=S(\Delta(\rho_{S}))$ lead us to the followings.
$E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\geq S(\sigma_{S})-S(\sigma_{SE})$
$=$ $S(\sigma_{S})-S(\rho_{S})-S(\rho_{E})$
$=$ $S(\sigma_{S})-S(\Delta(\rho_{S}))+S(\Delta(\rho_{S}))-S(\rho_{S})-S(\rho_{E})$
$=$ $C_{r}(\rho_{S})-C_{r}(\sigma_{S})-S(\rho_{E})$ (25)
Moving to the right hand side, from the definition of the relative entropy of entanglement: $E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\leq S(\sigma_{SE}||\sigma_{S}\otimes\sigma_{E})=I(S:E)=S(\sigma_{S})+S(\sigma_{E})-S(\sigma_{SE})=S(\sigma_{S})+S(\sigma_{E})-S(\rho_{S})-S(\rho_{E})$, since by definition $\sigma_{SE}=U\rho_{S}\otimes\rho_{E}U^{\dagger}$ so that $S(\sigma_{SE})=S(\rho_{S}\otimes\rho_{E})=S(\rho_{S})+S(\rho_{E})$. We also observe that $S(\sigma_{E})=S(\sum_{i}p_{i}\rho_{i})\leq S(\rho_{E})+H(i)$, which follows from the well known property of the von Neumann entropy: $S(\sum_{i}p_{i}\rho_{i})\leq\sum p_{i}S(\rho_{i})+H(i)$. Substituting this into the above inequality, we prove the statement. ∎
Lemma 4 fits the intuition that larger system-environment entanglement leads to more decoherence quantified by $C_{r}(\rho_{S})-C_{r}(\sigma_{S})$. In other words, decoherence in the system $D_{S}$ is lower bounded by generated entanglement.
4 Case study: The spin-boson model
In this Section, we illustrate our results on the spin-boson model, which is one of the most popular open systems models; see e.g. Schlosshauer (2007); Breuer and Petruccione (2007), and the references therein for an introduction. In its most used form, it describes a two-level central system, a qubit, interacting with a collection of bosonic modes (or harmonic oscillators), constituting the environment; for an example of higher dimensional central systems see e.g. Kiciński and Korbicz (2021). The dynamics of the model is given by the following Hamiltonian:
$H$ $=\sum_{\boldsymbol{k}}\omega_{\boldsymbol{k}}a^{\dagger}_{\boldsymbol{k}}a_{\boldsymbol{k}}+\sigma_{z}\otimes\sum_{\boldsymbol{k}}\left(g_{\boldsymbol{k}}a^{\dagger}_{\boldsymbol{k}}+g^{*}_{\boldsymbol{k}}a_{\boldsymbol{k}}\right),$ (26)
where for the purpose of this example, we have suppressed the self-Hamiltonian of the central system, or assumed it to be commuting with the interaction, as otherwise the evolution would not be of the pure dephasing form (4). Here, $a^{\dagger}_{\boldsymbol{k}}$, $a_{\boldsymbol{k}}$ are the annihilation and creation operators of the mode $\boldsymbol{k}$ of the environment, $\sigma_{z}$ is the $z$ axis Pauli matrix acting in the space of the central system and $g_{\boldsymbol{k}}$ are the coupling constants. The pointer basis here is the eigenbasis of $\sigma_{z}$, which follows immediately from the interaction term in (26).
We assume the environment to be thermal at temperature $T$, i.e. $\rho_{E}=e^{-\beta H_{E}}/Z$, where we choose the units so that $\beta=1/T$, and the central system to be initially in the following mixed state:
$\rho_{S}=(1-\alpha)|+\rangle\langle+|+\alpha\frac{\mathbbm{I}}{2}.$ (27)
It is a mixture of a state that is maximally coherent w.r.t. the pointer basis and the maximally mixed state $\mathbbm{I}/2$, and $\alpha$ is the mixedness parameter, $0\leq\alpha\leq 1$. Such a choice of the initial state will allow us to control the initial coherence in the pointer basis.
Since the central system is two-dimensional, we will use the bound (1). The initial coherence $C_{r}(\rho_{S})$ can be easily calculated and is given by the binary entropy:
$C_{r}(\rho_{S})=1-h\left(\frac{\alpha}{2}\right)$ (28)
Interestingly, the fidelity $B(\rho_{0}(t),\rho_{1}(t))$ can also be calculated analytically for the popular Ohmic spectral density:
$J(\omega)=\frac{\omega^{s}}{\Lambda^{s-1}}\bm{e}^{-\omega/\Lambda},$ (29)
where $\Lambda$ is the cutoff frequency and $s$ is the Ohmicity parameter. The value $s=3$ is a known threshold, distinguishing Markovian evolution for $s<3$ from non-Markovian for $s\geq 3$; see e.g. Addis et al. (2014). In Lampo et al. (2017) the analytical expression for $B$ was obtained for the case of the Ohmicity parameter $s>1$. It consists of two parts, vacuum, i.e. temperature-independent, and thermal:
$\ln B(t)=$ $\ln B_{vac}(t)+\ln B_{th}(t),$ (30)
$\frac{1}{2}\ln B_{vac}(t)=$ $-\wp(s-1)\left[1-\frac{\cos[(s-1)\arctan(\Lambda t)]}{(1+\Lambda^{2}t^{2})^{\frac{s-1}{2}}}\right],$ (31)
$\frac{1}{4}\ln B_{th}(t)=$ $\left(-\frac{T}{2\Lambda}\right)^{s-1}\bigg{\{}\Psi^{(s-2)}\left(1+\frac{T}{2\Lambda}\right)-\Psi^{(s-2)}\left(\frac{1}{2}+\frac{T}{2\Lambda}\right)$
[…]
where c.c. means complex conjugate of the last two terms in the parenthesis, $\wp(s)$ is the Euler gamma function, and $\Psi^{m}(z)$ is the polygamma function defined as:
$\Psi^{m}(z)=\sum_{k=0}^{\infty}\frac{(-1)^{m+1}m!}{(z+k)^{m+1}}.$ (33)
Figure 1: Lower bound (1) for the spin-boson model as a function of time for a different mixedness parameter $\alpha$ (cf. (27)) and different Ohmicity: a) $s=2$ corresponding to Markovian evolution; b) $s=3$ corresponding to non-Markovian evolution. The ratio of the thermal to the cutoff energy is set to $T/\Lambda=1$ and the time is measured in the cutoff units. The insert shows the peak and thus the non-monotonic behavior of the plots (with different vertical scale for each plot). On all plots the parts, where the bound becomes negative were replaced by the trivial bound equal to zero.
[…]
In Fig. 1, we keep fixed the ratio of the thermal to the cutoff energy, $T/\Lambda=1$, and plot (1) for different mixedness $\alpha$ and Ohmicity $s$ parameters. As expected from the (1), less mixed, i.e. more coherent, states with lower $\alpha$ lead to much more entanglement generation. More interestingly, there is a slight quantitative change of behavior between $s=2$ and $s=3$, corresponding to the transition between the Markovian and non-Markovian evolution respectively. While for $s=2$ the behavior is monotonic in time, for $s=3$ there is a broad peak around $\Lambda t=2$, regardless the value of $\alpha$ (depicted on the insert). This peak is more visible in Fig. 2, where we fix the mixedness $\alpha=1/2$ and vary the energy ratio $T/\Lambda$. The non-monotonic behavior of the entanglement bound for $s=3$ bears certain similarity to non-Markovianity measures, where a non-monotonic behavior of some marker function certifies information back-flow and hence the non-Markovianity of the evolution. Especially close seems the measure from Hall et al. (2014), where the non-Markovianity is detected via non-monotonicity of the decoherence rate, which in turn derives from system-environment correlations.
We also notice from Fig. 1 that the higher the $T/\Lambda$ ratio, the higher the entanglement bound, which means that if the cutoff frequency $\Lambda$ is fixed, higher temperatures guarantee stronger entanglement between the system and the environment. This in turn leads to a stronger decoherence and explains the well known fact why the system need to be kept cold in order to maintain the coherence.
5 Conclusions
We studied generation of entanglement between the system of interest and its environment for a broad class of so called pure dephasing evolution. We extended the previous results  Roszak and Cywiński (2015); Roszak (2018), obtained for pure initial states of the system, to the case of mixed states. We combined methods and results from coherence theory, state discrimination, and open systems evolution to obtain a lower bound on the generated entanglement, measured by the relative entropy.
Our second result describes another lower bound and upper bound for generated entanglement in terms of the difference of coherence in the system before and after the evolution. These bounds also can be seen as an extension of the result obtained in Streltsov et al. (2015). Although both the lower bound and upper bound contains the extra terms other than decoherence, we successfully prove that larger entanglement indicates more decoherence in pure dephasing evolution with arbitrary environmental state and vice versa.
Regarding the first result, its lower bound character is a price to pay for the generalization to mixed states. However, it certifies that at least the given amount of entanglement was created during the evolution. The first result is the strongest for the two-dimensional central system, where we obtained a compact formula. For arbitrary dimension, our bound still contains a certain optimization. As an example, we calculated analytically and studied numerically our bound for the popular spin-boson system, rediscovering the well-known traces of the non-Markovianity but in the entanglement dynamics. Obviously, more systems can be studied, especially that in some of them the fidelities has already been calculated Korbicz et al. (2014); Tuziemski et al. (2019); Kwiatkowski et al. (2021); Kiciński and Korbicz (2021)
Acknowledgements
We would like to thank K. Roszak for getting us interested in the topic and for fruitful discussions and comments. We acknowledge the financial support of the Polish National Science Foundation (NCN) through the grants Opus 2019/35/B/ST2/01896 and Quant-Era “Quantum Coherence Activation By Open Systems and Environments” QuCABOoSE 2023/05/Y/ST2/00139.